# Learning to Adapt: Bio-Inspired Gait Strategies for Versatile Quadruped Locomotion

### v1

@@ III-B Bio-inspired Gait Scheduler
The BGS primary output, $\bm{\beta}_{L}=[\bm{c}^{\text{ref}},\bm{p}_{x}^{\text{ref}},\bm{p}_{y}^{\text{ref}},\bm{p}_{z}^{\text{ref}}]\in{\mathbb{R}}^{\text{16}}$, defines the reference contact state of each foot, $\bm{c}^{\text{ref}}\in{\mathbb{B}}^{\text{4}}$, and their reference Cartesian position in the world frame $x$-axis, $\bm{p}_{x}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$, $y$-axis, $\bm{p}_{y}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$, and $z$-axis, $\bm{p}_{z}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$ which are calculated online using the Raibert heuristic [44] to account for the current state of the robot. An adjusted version of the BGS output, $\bm{\beta}_{G}$, is used for $\pi_{G}$ as not all the information in $\bm{\beta}_{L}$ is required. This has the form of $\bm{\beta}_{G}=[\bm{c}^{\text{ref}},\bm{p}_{z}^{\text{ref}},\Omega_{\text{stab}},\kappa]\in{\mathbb{R}}^{\text{10}}$ where $\Omega_{\text{stab}}\in\mathbb{R}$ characterises the inherent stability of a gait [27], and $\kappa\in\mathbb{B}$ is a logical flag to indicate a state of gait transition. We originally developed the BGS within [45] where the Froude number [27], $\Omega$, is used to trigger gait transition based exclusively on CoT which results in a set order of transitions. However, when applied to this work this method is not entirely suitable as now multiple biomechanics metrics and a set of auxiliary gaits need to be considered. One issue is that $\Omega$ values over $1$ are not compatible when calculating how many gait cycles, $C$, should a transition occur over. With this work investigating higher velocities than in [45], this has been resolved through calculating $C$ through
@@
where $\Omega_{\text{stab}}^{\text{next}}$ is the $\Omega_{\text{stab}}$ of the gait that’s being transitioned to. In essence, $f$ of the current and next gait dictates the harshness of the transition. This behaviour is also reflected in animal gait transitions, where transitioning from running (higher $f$) to trotting (lower $f$) the transition is slower compared to the opposite scenario [46]. Overall, this augmented version of the BGS can achieve transition between any designed gait, while considering the inherent stability of the transition. Complete details of how $\bm{c}^{\text{ref}}$ is generated for each gait can be found in Supplementary Information C (Appendix A-C).
@@ III-C Policy Training
in which $\gamma\in[0,1)$ is the discount factor, $\xi$ is a finite-horizon trajectory dependent on $\pi$ with length $N$, $p(\xi|\pi)$ is the likelihood of $\xi$, and $r$ is the reward function. The proximal policy optimization (PPO) algorithm [47] is used to train the locomotion policy and the hyperparameters used are detailed in Supplementary Information D (Appendix A-D). As discussed in Section I-A, we estimate the state of the robot during training using an SE. Hence, in terms of applying state feedback noise for domain randomisation to improved sim-to-real transfer we only need to implement this on the input sensor data vector of the SE, $\bm{\sigma}=[\bm{\omega_{B}},\dot{\bm{v}}_{B},\bm{q},\dot{\bm{q}},\bm{\tau},\bm{f}_{\text{grf}}]$. This vector includes base angular velocity, $\bm{\omega_{B}}\in{\mathbb{R}}^{\text{3}}$, base linear acceleration, $\dot{\bm{v}}_{B}\in{\mathbb{R}}^{\text{3}}$, joint positions, $\bm{q}$, joint velocities, $\dot{\bm{q}}$, joint torques, $\bm{\tau}$, and foot ground reaction forces, $\bm{f}_{\text{grf}}\in{\mathbb{R}}^{\text{4}}$. As the initial state of the robot and its performance can never be guaranteed during real-world deployment, we also randomise the initial configuration of the robot, the mass of the robot’s base, $K_{p}$ and $K_{d}$. Additionally, to ensure that a rich variation of $\bm{U}^{\text{cmd}}$ is experienced during training randomly sampled gaits, velocity commands and velocity change durations (to achieve random acceleration) are implemented during training. For all details regarding the noise and sampling used within training please refer to Supplementary Information E (Appendix A-E). The environment itself is constructed using RaiSim [48], as the vectorized environment setup allows for efficient training of policies. Additionally, the observation normalisation functionality offered by RaiSim is also used for improved training.
@@ III-D Locomotion Policy
The goal of the locomotion policy $\pi_{L}$ is to realise the input $\bm{U}^{\text{cmd}}$ while exhibiting stable and versatile behaviour. As such, $\pi_{L}$ is trained to generate the action, $\bm{q}^{*}$, from an input observation, $\bm{o}_{L}=[\bm{\beta}_{L},\bm{s},\bm{v}_{B}^{\text{cmd}}]\in{\mathbb{R}}^{\text{69}}$, where $\bm{v}_{B}^{\text{cmd}}=[v_{x}^{\text{cmd}},v_{y}^{\text{cmd}},\omega_{z}^{\text{cmd}}]\in{\mathbb{R}}^{\text{3}}$ is the high-level velocity command of the robot’s base within $\bm{U}^{\text{cmd}}$, as outlined in Fig. 1. $\bm{s}$ is generated from the output of the SE and is defined as $\bm{s}=[\bm{\alpha}\bm{R}_{B}^{T},\bm{q},\bm{\omega}_{B},\dot{\bm{q}},\bm{v}_{B},z_{B},\bm{\tau},\bm{c}]\in{\mathbb{R}}^{\text{50}}$, where $\bm{\alpha}=[0,0,1]^{T}$ is used to select the vertical $z$-axis, $\bm{\omega}_{B}\in{\mathbb{R}}^{\text{3}}$ is the base angular velocity, $\bm{v}_{B}\in{\mathbb{R}}^{\text{3}}$ is the base linear velocity, $z_{B}$ is the current base height, and $\bm{c}\in{\mathbb{B}}^{\text{4}}$ is the contact state of the feet. The locomotion reward function, $r_{L}$, is formulated so that the the output of the policy can realise the reference gait patterns and velocity commands stably, smoothly and accurately,
@@
where $\bm{p}_{i}\in{\mathbb{R}}^{\text{3}}$ is the velocity of the $i$-th foot scheduled to be in stance, $F$ is the number of stance feet, $\bm{\omega}_{B}=[\omega_{x},\omega_{y}]\in{\mathbb{R}}^{\text{2}}$ where $\omega_{x}$ and $\omega_{y}$ are roll and pitch base velocities respectively, $\bm{R}_{B}^{\text{des}}\in SO(\text{3})$ is the desired base orientation, $z_{B}^{\text{nom}}$ is the nominal base heights respectively, and $\bm{q}_{\text{hip}}\in{\mathbb{R}}^{\text{4}}$ is the hip angular joint positions.
@@ III-E Biomechanics Gait Transition Metrics
As the set of biomechanics metrics applied in this work were originally designed for analysing animal locomotion, several adjustments to how they are calculated needed to be implemented; for example, energy consumption in animals is often measured through the rate of consumption of $\text{O}_{2}$, hence unsuitable for the application of robotics. Additionally, as robots provide a wide array of feedback data some of the metrics have also been augmented to better reflect the characteristic that these biomechanics metrics are attempting to characterise. That being said, for Fig. 5(a) only the original biomechanics metrics are applied to allow for direct comparison between robot and animal data.
@@ III-E4 Stability
The best indication of stability in animals is their stride duration CV. This metric characterises periodicity, which is a primary indication of stable locomotion [31]. However, to accurately calculate this, the mean and standard deviation of the stride duration needs to be taken over an extended period of time for appropriate data generation. This is sufficient for undertaking analysis similar to that presented in Fig. 5(a), but this presents an issue when it comes to analysing the performance of the proposed control framework as it is common for multiple speed commands being used within the duration of one stride. Hence, to overcome this limitation we instead use $c_{\text{avg}}^{\text{err}}=|\bm{c}^{\text{err}}|/4$, which can be measured every time step rather than just each foot touchdown event; the gait references generated by the BGS have a constant and periodic stride duration, therefore an accurate tracking of this reference would in turn indicate high periodicity, which is further supported by the correlation between the two metrics within Fig. 5(b).
@@ III-F Gait Selection Policy
To achieve optimal gait selection for a given state, we leverage the biomechanics metrics within the reward function of $\pi_{G}$, $r_{G}$. For the different variations of $\pi_{G}$ used within Fig. 5(a), each policy’s reward function only features the metric that its focusing on within $r_{G}$ but $\pi_{G}^{\text{uni}}$ unifies all metrics hence uses the full form of $r_{G}$ with all metrics. Additionally, as the biomechanics metrics all describe a characteristic that animals try to minimise through changing gaits, they can be directly applied within $r_{G}$ with some normalisation where appropriate. The full form of $r_{G}$ is
@@
in which $r_{v^{\text{cmd}}}$ is taken from (7), and $r_{\text{stand}}$ is set to $-10$ if a stand gait is used when $|\bm{v}_{B}^{\text{cmd}}|>0$ or not used when $|\bm{v}_{B}^{\text{cmd}}|=0$. For $r_{\text{smooth}}$, the reward aims to penalise unnecessary changes in $\Gamma^{*}$ to remove rapid gait changes when two gaits could achieve similar metric minimisation for a given task and state. As such, if there is a gait change between time steps it is calculated as $r_{\text{smooth}}=-\psi(\text{CoT}+\tau_{\%}+c_{\text{avg}}^{\text{err}}+W_{\text{ext}})$ otherwise it is set to 00. To generate $\Gamma^{*}$, $\pi_{G}$ takes in input observation vector $\bm{o}_{G}=[\bm{s},\bm{\beta}_{G},\bm{v}_{B}^{\text{cmd}},\dot{\bm{v}}_{B}^{\text{cmd}},\Gamma^{*}_{t-1}]\in{\mathbb{R}}^{\text{66}}$ in which $\Gamma^{*}_{t-1}$ is the previous output action to aid in action smoothing. Appropriate selection of the data provided to $\pi_{G}$ is critical in order to achieve targeted minimisation of the biomechanics metrics. As such, the inclusion of $\bm{s}$ coupled with $\bm{c}^{\text{ref}}$, $\bm{p}_{z}^{\text{ref}}$, $\bm{v}_{B}^{\text{cmd}}$, $\dot{\bm{v}}_{B}^{\text{cmd}}$ and $\Omega_{\text{stab}}$ informs the policy of its current and demanded stability, while the terms $\bm{\tau}$ and $\dot{\bm{q}}$ within $\bm{s}$ capture the power consumption of the robot and the forces to which it is subjected.
Appendix A Supplementary Information
A-A Preservation of Biomechanics Metrics on Grass
Figure A.1: Deployment of $\pi_{L}^{\text{bio}}$ and $\pi_{G}^{\text{uni}}$ on smooth low-friction floor and uneven grassy terrain, with the black triangles indicating a point of failure.
@@ A-A Preservation of Biomechanics Metrics on Grass
When deploying $\pi_{L}^{\text{bio}}$ with $\pi_{G}^{\text{uni}}$ on a smooth low-friction floor and uneven grassy terrain $\pi_{G}^{\text{uni}}$ can utilise the same behaviour animals demonstrate in minimising CoT; comparing the performance of maintaining a static gait of both trot and run gaits within Fig. A.1, $\pi_{G}^{\text{uni}}$ can select the most energy efficient gait and even prevent failure as the trot gait cannot maintain stability for the entire experiment. Additionally, the distribution of gait usage between trot and run gaits varies between the two terrains. This is due to the uneven grassy terrain causing reduced stability of the robot when trotting, as indicated by $c_{\text{avg}}^{\text{err}}$ in Fig. A.1, which in response $\pi_{G}^{\text{uni}}$ exhibits increased modulation between trotting and running to preserve stability. This preservation of stability while minimising CoT is achieved through the resultant modulation of the stride frequency as gaits with higher stride frequency offer improved stability [31] and efficiency on rough terrain [29].
A-B Stability Recovery from External Perturbations
@@
An additional stability recovery experiment was completed where the robot was subjected to repeated external perturbations and in response utilised the utility gaits to recover, as presented in Fig. A.2, which in turn reflects the findings discussed in Sections I-D and II.
@@
Figure A.2: Enacting perturbations upon the legs of the robot to investigate how our framework can reject this instability.
A-C Generation Gait Contact References within the BGS
TABLE A.1: Gait design parameters.
@@ A-C Generation Gait Contact References within the BGS
In accordance to our work in [45], to generate $\bm{c}^{\text{ref}}$, a set of phase variables, $\phi_{i}\in[0,1)$, for each leg $\phi_{1},\ldots,\phi_{4}$, are used to determine the progress along a gait pattern, and duty factor, $d_{i}\in[0,1)$, sets the percentage of the phase that each leg is in stance. These encoded parameters for each gait are presented within Table A.1. When $\phi_{i}=d_{i}$, the contact state of the $i$-th leg switches to swing. The phase of the $i$-th leg is calculated as
$\phi_{i}=\frac{t-t_{i,0}}{T},$ (16)
@@
A-D PPO Hyperparameters
All PPO hyperparameters used within this work are detailed within Table A.2.
TABLE A.2: PPO Hyperparameters.
@@ A-D PPO Hyperparameters
A-E Noise and Sampling Distributions Used During Training
TABLE A.3: Noise and Sampling Distributions.

### v2

@@ III-B Bio-inspired Gait Scheduler
The BGS primary output, $\bm{\beta}_{L}=[\bm{c}^{\text{ref}},\bm{p}_{x}^{\text{ref}},\bm{p}_{y}^{\text{ref}},\bm{p}_{z}^{\text{ref}}]\in{\mathbb{R}}^{\text{16}}$, defines the reference contact state of each foot, $\bm{c}^{\text{ref}}\in{\mathbb{B}}^{\text{4}}$, and their reference Cartesian position in the world frame $x$-axis, $\bm{p}_{x}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$, $y$-axis, $\bm{p}_{y}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$, and $z$-axis, $\bm{p}_{z}^{\text{ref}}\in{\mathbb{R}}^{\text{4}}$ which are calculated online using the Raibert heuristic [48] to account for the current state of the robot. Throughout this paper, the limits enforced on the generation of $\bm{p}_{x}^{\text{ref}}$, $\bm{p}_{y}^{\text{ref}}$ and $\bm{p}_{z}^{\text{ref}}$ are $0.3$ m, $0.2$ m and $0.1$ m respectively from the nominal local foot position. An adjusted version of the BGS output, $\bm{\beta}_{G}$, is used for $\pi_{G}$ as not all the information in $\bm{\beta}_{L}$ is required. This has the form of $\bm{\beta}_{G}=[\bm{c}^{\text{ref}},\bm{p}_{z}^{\text{ref}},\Omega_{\text{stab}},\kappa]\in{\mathbb{R}}^{\text{10}}$ where $\Omega_{\text{stab}}\in\mathbb{R}$ characterises the inherent stability of a gait [5], and $\kappa\in\mathbb{B}$ is a logical flag to indicate a state of gait transition. We originally developed the BGS within [49] where the Froude number [5], $\Omega$, is used to trigger gait transition based exclusively on CoT which results in a set order of transitions. However, when applied to this work this method is not entirely suitable as now multiple biomechanics metrics and a set of auxiliary gaits need to be considered. One issue is that $\Omega$ values over $1$ are not compatible when calculating how many gait cycles, $C$, should a transition occur over. With this work investigating higher velocities than in [49], this has been resolved through calculating $C$ through
@@
where $\Omega_{\text{stab}}^{\text{next}}$ is the $\Omega_{\text{stab}}$ of the gait that’s being transitioned to. In essence, $f$ of the current and next gait dictates the harshness of the transition. This behaviour is also reflected in animal gait transitions, where transitioning from running (higher $f$) to trotting (lower $f$) the transition is slower compared to the opposite scenario [50]. Overall, this augmented version of the BGS can achieve transition between any designed gait, while considering the inherent stability of the transition. Complete details of how $\bm{c}^{\text{ref}}$ is generated for each gait can be found in Supplementary Section 5.
@@ III-C Policy Training
in which $\gamma\in[0,1)$ is the discount factor, $\xi$ is a finite-horizon trajectory dependent on $\pi$ with length $N$, $p(\xi|\pi)$ is the likelihood of $\xi$, and $r$ is the reward function. The proximal policy optimization (PPO) algorithm [51] is used to train all policies and the hyperparameters used are detailed in Supplementary Section 6, which were selecting through the standard method of parameter tuning. As discussed in Section I-A, we estimate the state of the robot during training using an SE. Hence, in terms of applying state feedback noise for domain randomisation to improved sim-to-real transfer we only need to implement this on the input sensor data vector of the SE, $\bm{\sigma}=[\bm{\omega_{B}},\dot{\bm{v}}_{B},\bm{q},\dot{\bm{q}},\bm{\tau},\bm{f}_{\text{grf}}]$. This vector includes base angular velocity, $\bm{\omega_{B}}\in{\mathbb{R}}^{\text{3}}$, base linear acceleration, $\dot{\bm{v}}_{B}\in{\mathbb{R}}^{\text{3}}$, joint positions, $\bm{q}$, joint velocities, $\dot{\bm{q}}$, joint torques, $\bm{\tau}$, and foot ground reaction forces, $\bm{f}_{\text{grf}}\in{\mathbb{R}}^{\text{4}}$. As the initial state of the robot and its performance can never be guaranteed during real-world deployment, we also randomise the initial configuration of the robot, the mass of the robot’s base, $K_{p}$ and $K_{d}$. Additionally, to ensure that a rich variation of $\bm{U}^{\text{cmd}}$ is experienced during training randomly sampled gaits, velocity commands and velocity change durations (to achieve random acceleration) are implemented during training. For all details regarding the noise and sampling used within training please refer to Supplementary Section 7. Although sim-to-real transfer can pose a significant challenge when training DRL policies, we have found that through using domain randomisation, realistic and diverse velocity commands, and generating all robot state observations from the state estimator our framework is able to achieve zero-shot traversal in all experiments and environments shown in Figures 2 and 6, hence demonstrating our methods sufficiently bridge the gap between simulation and the real world. The environment itself is constructed using RaiSim [52], as the vectorized environment setup allows for efficient training of policies. Additionally, the observation normalisation functionality offered by RaiSim is also used for improved training.
@@ III-D Locomotion Policy
The goal of the locomotion policy $\pi_{L}$ is to realise the input $\bm{U}^{\text{cmd}}$ while exhibiting stable and versatile behaviour. As such, $\pi_{L}$ is trained to generate the action, $\bm{q}^{*}$, from an input observation, $\bm{o}_{L}=[\bm{\beta}_{L},\bm{s},\bm{v}_{B}^{\text{cmd}}]\in{\mathbb{R}}^{\text{69}}$, where $\bm{v}_{B}^{\text{cmd}}=[v_{x}^{\text{cmd}},v_{y}^{\text{cmd}},\omega_{z}^{\text{cmd}}]\in{\mathbb{R}}^{\text{3}}$ is the high-level velocity command of the robot’s base within $\bm{U}^{\text{cmd}}$, as outlined in Figure 1. $\bm{s}$ is generated from the output of the SE and is defined as $\bm{s}=[\bm{\alpha}\bm{R}_{B}^{T},\bm{q},\bm{\omega}_{B},\dot{\bm{q}},\bm{v}_{B},z_{B},\bm{\tau},\bm{c}]\in{\mathbb{R}}^{\text{50}}$, where $\bm{\alpha}=[0,0,1]^{T}$ is used to select the vertical $z$-axis, $\bm{\omega}_{B}\in{\mathbb{R}}^{\text{3}}$ is the base angular velocity, $\bm{v}_{B}\in{\mathbb{R}}^{\text{3}}$ is the base linear velocity, $z_{B}$ is the current base height, and $\bm{c}\in{\mathbb{B}}^{\text{4}}$ is the contact state of the feet. The locomotion reward function, $r_{L}$, is formulated so that the the output of the policy can realise the reference gait patterns and velocity commands stably, smoothly and accurately,
@@
where $\bm{p}_{i}\in{\mathbb{R}}^{\text{3}}$ is the velocity of the $i$-th foot scheduled to be in stance, $F$ is the number of stance feet, $\bm{\omega}_{B}=[\omega_{x},\omega_{y}]\in{\mathbb{R}}^{\text{2}}$ where $\omega_{x}$ and $\omega_{y}$ are roll and pitch base velocities respectively, $\bm{R}_{B}^{\text{des}}\in SO(\text{3})$ is the desired base orientation, $z_{B}^{\text{nom}}$ is the nominal base heights respectively, and $\bm{q}_{\text{hip}}\in{\mathbb{R}}^{\text{4}}$ is the hip angular joint positions. Overall, this reward function enables deployment of all targeted gaits with rapid transitions between them, even at high speeds, as demonstrated in Figures 3, 4 and 6.
@@ III-E Biomechanics Gait Transition Metrics
As the set of biomechanics metrics applied in this work were originally designed for analysing animal locomotion, several adjustments to how they are calculated needed to be implemented; for example, energy consumption in animals is often measured through the rate of consumption of $\text{O}_{2}$, hence unsuitable for the application of robotics. Additionally, as robots provide a wide array of feedback data some of the metrics have also been augmented to better reflect the characteristic that these biomechanics metrics are attempting to characterise. That being said, for Figure 5(a) only the original biomechanics metrics are applied to allow for direct comparison between robot and animal data.
@@ III-E4 Stability
The best indication of stability in animals is their stride duration CV. This metric characterises periodicity, which is a primary indication of stable locomotion [35]. However, to accurately calculate this, the mean and standard deviation of the stride duration needs to be taken over an extended period of time for appropriate data generation. This is sufficient for undertaking analysis similar to that presented in Figure 5(a), but this presents an issue when it comes to analysing the performance of the proposed control framework as it is common for multiple speed commands being used within the duration of one stride. Hence, to overcome this limitation we instead use $c_{\text{avg}}^{\text{err}}=|\bm{c}^{\text{err}}|/4$, which can be measured every time step rather than just each foot touchdown event; the gait references generated by the BGS have a constant and periodic stride duration, therefore an accurate tracking of this reference would in turn indicate high periodicity, which is further supported by the correlation between the two metrics within Figure 5(b).
@@ III-F Gait Selection Policy
To achieve optimal gait selection for a given state, we leverage the biomechanics metrics within the reward function of $\pi_{G}$, $r_{G}$. For the different variations of $\pi_{G}$ used within Figure 5(a), each policy’s reward function only features the metric that its focusing on within $r_{G}$ but $\pi_{G}^{\text{uni}}$ unifies all metrics hence uses the full form of $r_{G}$ with all metrics. Additionally, as the biomechanics metrics all describe a characteristic that animals try to minimise through changing gaits, they can be directly applied within $r_{G}$ with some normalisation where appropriate. The full form of $r_{G}$ is
@@
in which $r_{v^{\text{cmd}}}$ is taken from (7), and $r_{\text{stand}}$ is set to $-10$ if a stand gait is used when $|\bm{v}_{B}^{\text{cmd}}|>0$ or not used when $|\bm{v}_{B}^{\text{cmd}}|=0$. For $r_{\text{smooth}}$, the reward aims to penalise unnecessary changes in $\Gamma^{*}$ to remove rapid gait changes when two gaits could achieve similar metric minimisation for a given task and state. As such, if there is a gait change between time steps it is calculated as $r_{\text{smooth}}=-\psi(\text{CoT}+\tau_{\%}+c_{\text{avg}}^{\text{err}}+W_{\text{ext}})$ otherwise it is set to 00. To generate $\Gamma^{*}$, $\pi_{G}$ takes in input observation vector $\bm{o}_{G}=[\bm{s},\bm{\beta}_{G},\bm{v}_{B}^{\text{cmd}},\dot{\bm{v}}_{B}^{\text{cmd}},\Gamma^{*}_{t-1}]\in{\mathbb{R}}^{\text{66}}$ in which $\Gamma^{*}_{t-1}$ is the previous output action to aid in action smoothing. Appropriate selection of the data provided to $\pi_{G}$ is critical in order to achieve targeted minimisation of the biomechanics metrics. As such, the inclusion of $\bm{s}$ coupled with $\bm{c}^{\text{ref}}$, $\bm{p}_{z}^{\text{ref}}$, $\bm{v}_{B}^{\text{cmd}}$, $\dot{\bm{v}}_{B}^{\text{cmd}}$ and $\Omega_{\text{stab}}$ informs the policy of its current and demanded stability, while the terms $\bm{\tau}$ and $\dot{\bm{q}}$ within $\bm{s}$ capture the power consumption of the robot and the forces to which it is subjected. Overall, through the formulation of the biomechanics metrics within this reward function, we are not only able to full investigate the effects in gait selection when minimising each metric, but also instil the intrinsics of animal gait transition strategies within a DRL gait selection policy, as detailed in Figure 5.
Data Availability
The simulation and hardware experimental data used to generate the figures within this paper are available at https://github.com/ihcr/learning_to_adapt [54]. The animal data within the repository was extracted from [35, 46, 33], [37] and [33].
Code Availability
The code and demos of our framework is available on GitHub at https://github.com/ihcr/learning_to_adapt [54].
S1 Classification of Froude-characterised and Froude-free Locomotion
Terrestrial locomotion can take on many forms but when it comes to animal quadruped locomotion two main groups can be identified, locomotion that can be characterised by the Froude number [5] and locomotion that cannot:
  • Froude-characterised locomotion
    • This includes walking, running, hopping, ambling, pronking, pacing, bounding, limping, cantering, galloping and other rhythmic gaits.
    • Can account for $\approx 70-90\%$ of the daily locomotion utilised by terrestrial animals [13, 14].
    • Gait transition need to occur at both low and high velocity magnitudes, hence requiring the ability to rapidly change gait under dynamic conditions.
    • Features a desired velocity that has not normal component to the ground plane.
  • Froude-free locomotion
    • This includes climbing, jumping, crouching and other non-rhythmic gaits or motion skills.
    • Can account for $\approx 10-30\%$ of the daily locomotion utilised by terrestrial animals [13, 14].
    • Gait and skill transitions occur at low velocity magnitudes due to the extreme differences in motion.
    • Features a desired velocity that has normal components to the ground plane.
    • Breaks the assumptions that are used in the formulation of the Froude number due to large variations in the height of the base relative to the ground.
With Froude-characterised locomotion being the primary locomotion utilised by animals to traverse natural terrain, this exemplifies the importance of quadruped robots mastering this skill for optimal performance in the majority of scenarios. Consequently, it is this highly adaptive gait-based locomotion and transition strategies seen in animals that we aimed to instil within our locomotion framework for optimal Froude-characterised locomotion.
S2 Preservation of Biomechanics Metrics on Grass
@@
When deploying $\pi_{L}^{\text{bio}}$ with $\pi_{G}^{\text{uni}}$ on a smooth low-friction floor and uneven grassy terrain $\pi_{G}^{\text{uni}}$ can utilise the same behaviour animals demonstrate in minimising CoT; comparing the performance of maintaining a static gait of both trot and run gaits within Supplementary Figure 1, $\pi_{G}^{\text{uni}}$ can select the most energy efficient gait and even prevent failure as the trot gait cannot maintain stability for the entire experiment. Additionally, the distribution of gait usage between trot and run gaits varies between the two terrains. This is due to the uneven grassy terrain causing reduced stability of the robot when trotting, as indicated by $c_{\text{avg}}^{\text{err}}$ in Supplementary Figure 1, which in response $\pi_{G}^{\text{uni}}$ exhibits increased modulation between trotting and running to preserve stability. This preservation of stability while minimising CoT is achieved through the resultant modulation of the stride frequency as gaits with higher stride frequency offer improved stability [35] and efficiency on rough terrain [33]. Overall, on smooth low-friction floor $\pi_{G}^{\text{uni}}$ is able to achieve a $15\%$ and $24\%$ reduction in CoT compared to trotting and running gaits and a reduction of $18\%$ and $30\%$ on uneven grass.
Figure S1: Deployment of $\pi_{L}^{\text{bio}}$ and $\pi_{G}^{\text{uni}}$ on smooth low-friction floor and uneven grassy terrain, with the black triangles indicating a point of failure.
S3 Stability Recovery from External Perturbations
@@
An additional stability recovery experiment was completed where the robot was subjected to repeated external perturbations and in response utilised the auxiliary gaits to recover, as presented in Supplementary Figure 2, which in turn reflects the findings discussed in Sections I-D and II.
@@
Figure S2: Enacting perturbations upon the legs of the robot to investigate how our framework can reject this instability.
S4 Comparison to Other Froude-characterised Locomotion Frameworks
S4.1 Performance and Characteristics Comparison
To benchmark the performance of our framework, we have compiled all available performance metrics and characteristics of other similar frameworks in Supplementary Table 1, with the most relevant ones being:
  • DeepTransition [27]
  • WalkTheseWays [23]
  • RL+MPC [21]
  • Phase-guided [22]
  • EnergyLocomotion [24]
Comparing our framework to the others that take a bio-inspired approach, DeepTransition and EnergyLocomotion, it has been deployed in a much wider variety of terrains. In respect of DeepTransition, this disparity could be due to their limited number of deployable gaits and their method of utilising a bio-inspired approach at the low-level, which as discussed in the main text, would limit its performance even though both our methods and theirs take inspartation from stability and efficiency animal gait strategies. In the case of EnergyLocomotion, as only one energy efficiency animal gait strategy was implemented, only three gaits emerged which in turn results in limited adaptability in comparison to our framework.
TABLE S1: Comparison between Froude-characterised Locomotion Frameworks
Phase-guided WalkTheseWays RL+MPC DeepTransition EnergyLocomotion Ours
Framework CPG-RL RL RL-MPC CPG-RL-IK RL RL
Type
Amble, Pronk, Trot, Pace Amble, Pronk Amble, Run
Gaits Pace, Trot, Pace, Trot, Pronk, Bound Trot, Pronk Amble, Trot, Pronk, Bound
Bound Bound Gallop Bound Limp, Amble, Hop
Level of Low-Level High-level for High-level for
Bio-inspired Stability and Efficiency Stability and
Implementation Efficiency Efficiency
Optimal Gait
Selection
Arbitrary Gait
Transitions
Hardware Black Panther Go1 Go1 A1 A1 A1
Wood-chip, Deep Rocks,
Large Step, Concrete Slabs,
Terrains Grass Sand, Steps, Grass,Tarmac, Flat Terrain Cardboard Sheets Tarmac, Loose Timber,
Traversed Grass, Wood-chip Wooden Boards with Large Gaps Grass, Rocks, Bush, Overgrown Roots, Leaves,
Wooden Planks Grass, Balanced Timber,
Low-friction Ramp
Additionally, both these frameworks cannot achieve arbitrary gait transitions as their gait selection strategies are embedded within their locomotion method, while our framework is able to achieve this due to $\pi_{L}^{\text{bio}}$ and $\pi_{G}^{\text{uni}}$ being independent policies which in turn promotes its suitability to be combined with Froude-free locomotion frameworks, as will be discussed in the next section. In the comparison between our framework and those who haven’t implemented neither optimal gait selection nor implemented any animal gait strategies (WalkTheseWays, RL+MPC and Phase-guided), with our framework being able to successfully traversing a much more diverse set of terrains and realise more types of gaits, this supports the notion that our method provides considerable improvements in adaptability and proficiency in Froude-characterised locomotion.
S4.2 Comparison Between Froude-characterised to Froude-free Focused Frameworks
Comparison between Froude-characterised focused to Froude-free focused locomotion frameworks is somewhat impractical. Primarily, Froude-free locomotion frameworks measure success through failure rate over challenging terrain and stability, with no concern or analysis on efficiency. Additionally, these frameworks are only able to realise a walking or running gait, with other Froude-characterised gaits and optimal gait selection being outside their operational scope. On the other hand, our framework, that focuses on Froude-characterised locomotion, determines its success based on command tracking error, rapid but stable gait transitions, efficiency and adaptability. Consequently, direct comparison between these two locomotion groups is not constructive as they are inherently different in terms of the challenges they are designed to overcome. However, an important and valuable direction for further research would be investigating how to effectively combine both framework types.
For example, although the Froude-free locomotion could potentially overcome the terrains featured in Figure 6 of the main text, they would do so without minimising energy consumption, with increased energy expenditure likely to be caused by excessive motion; as we have proven that the terrains in Figure 6 are traversable through just switching Froude-characterised gaits, use of Froude-free skills like climbing, jumping and sure-footedness would incur unreasonable energy costs. However, as we explicitly focus on optimal Froude-characterised locomotion in our work, through instilling animal gait strategies, our framework is not only able to traverse these terrains while preserving efficiency through optimal gait selection, but also recover from critically unstable states without requiring the use of inefficient Froude-free locomotion through rapid gait transitions to auxiliary gaits and then back to a nominal gait in a matter of milliseconds, as exhibited in Figure 6.
S5 Generation Gait Contact References within the BGS
TABLE S2: Gait design parameters.
@@
In accordance to our work in [49], to generate $\bm{c}^{\text{ref}}$, a set of phase variables, $\phi_{i}\in[0,1)$, for each leg $\phi_{1},\ldots,\phi_{4}$, are used to determine the progress along a gait pattern, and duty factor, $d_{i}\in[0,1)$, sets the percentage of the phase that each leg is in stance. These encoded parameters for each gait are presented within Supplementary Table 2. When $\phi_{i}=d_{i}$, the contact state of the $i$-th leg switches to swing. The phase of the $i$-th leg is calculated as
$\phi_{i}=\frac{t-t_{i,0}}{T},$ (S1)
@@
S6 PPO Hyperparameters
All PPO hyperparameters used within this work are detailed within Supplementary Table 3.
TABLE S3: PPO Hyperparameters.
@@
S7 Noise and Sampling Distributions Used During Training
In the effort of improving sim-to-real transfer, domain randomisation was utilised through randomly sampled noise from either uniform or normal distributions to all parameters within $\bm{\sigma}$, the initial configuration of the robot in each episode, $\bm{q}_{\text{init}}$, the mass of the robot’s base, $m_{B}$, teh fiction coefficient between the robot’s foot and the ground, $\mu_{\text{fric}}$, $K_{p}$ and $K_{p}$. The details of this is presented in Supplementary Table 4. Additionally, to ensure that a rich variation of $\bm{U}^{\text{cmd}}$ is experienced during training, randomly sampled gaits, velocity commands and velocity change durations (to achieve random acceleration), $t_{\text{acc}}$, are implemented during training. These are within defined maximum, $t_{\text{acc}}^{\text{max}}=0.5\text{s}$, and minimum, $t_{\text{acc}}^{\text{min}}=0\text{s}$, acceleration durations. Further sampling details are also presented in Supplementary Table 4.
TABLE S4: Noise and Sampling Distributions.